OpenStudy (anonymous):

Which value of x makes the equation true? x – 5.1 = –7.6 x =

OpenStudy (anonymous):

Try doing this on your own then i'll see if ur right :D Do what I did on the last question

OpenStudy (igreen):

Okay, we want to add 5.1 to both sides, can you do that?

OpenStudy (anonymous):

i think its 10.1

OpenStudy (anonymous):

No

OpenStudy (anonymous):

awwww

OpenStudy (igreen):

x - 5.1 + 5.1 = -7.6 + 5.1 -5.1 + 5.1 = ? -7.6 + 5.1 = ?

OpenStudy (anonymous):

oh ok i think its -2.5

OpenStudy (igreen):

You got it.

OpenStudy (igreen):

x - 5.1 + 5.1 = -7.6 + 5.1 x - 0 = -2.5 x = -2.5

OpenStudy (anonymous):

Good job @WOLFPACK13

OpenStudy (anonymous):

yayyyyyyy thanks guys ur the best

OpenStudy (anonymous):

byeeeeeeee